# Invariant Reduction for Partial Differential Equations. III: Poisson Brackets

### v2

@@ Abstract
We show that, under suitable conditions, finite-dimensional systems describing invariant solutions of partial differential equations (PDEs) inherit local Hamiltonian operators through the mechanism of invariant reduction, which applies uniformly to point, contact, and higher symmetries. The inherited operators endow the reduced systems with Poisson bivectors that relate constants of motion to symmetries. Applying the same mechanism to invariant conservation laws, we further show that the induced Poisson brackets agree with those of the original systems, up to sign. This is illustrated by two examples in which the inherited Poisson brackets and inherited constants of motion yield integrability of the reduced systems.
@@ 1 Introduction
Our mechanism of invariant reduction is related to methods introduced in [5] and [6]. As formulated in those works, these methods are not applicable to higher symmetries, but they do allow one to consider multi-reduction in a more general setting. Moreover, following the ideas of this paper, a slight extension of the method of [5] to the reduction of the invariant part of the zero page of Vinogradov’s $\mathcal{C}$-spectral sequence for a given Hamiltonian equation may allow Poisson bracket reductions under suitable, in some cases non-solvable, multi-dimensional Lie algebras of point or contact symmetries.
@@
Aside from this, there is a substantial literature on related questions concerning finite-dimensional symmetry reductions and the inheritance of Poisson brackets for evolution systems (see, e.g., [10, 11, 12, 13, 14, 15, 16, 17, 18]). Our approach is distinguished by its generality and systematic nature: assuming suitable regularity conditions, it applies to $\ell$-normal systems of partial differential equations, with (1+1)-dimensional evolution systems included as a particular case. We also emphasize that our reduction mechanism is global and geometric. In particular, it does not rely on local coordinates, auxiliary variables for reduced systems, or the existence of Lax pairs. While some aspects of symmetry reductions and Poisson bracket inheritance have been studied previously, the generality and formulation of our approach—including Theorem 7.1—appear to be novel.
@@ 2.1 Jets
Here $\langle\cdot,\cdot\rangle$ denotes the natural pairing between a module and its adjoint. Note that $\mathrm{E}(L)$ can be identified with the differential form that takes the following expression in adapted coordinates:
@@ 2.2 Differential equations
Regularity assumptions. We consider only systems that satisfy the following conditions.
  1. 1.
    For any point $r\in\{F=0\}\subset J^{k}(\pi)$, the differentials $dF^{i}_{r}$ of the components $F^{i}$ are linearly independent.
  2. 2.
    A function $f\in\mathcal{F}(\pi)$ vanishes on $\mathcal{E}$ if and only if there is a differential operator $\Delta\colon P(\pi)\to\mathcal{F}(\pi)$ of the form $\Delta_{i}^{\alpha}D_{\alpha}$ ($\mathcal{C}$-differential operator) such that $f=\Delta(F)$. Here, for some integer $k$, the components $\Delta_{i}^{\alpha}$ with $|\alpha|\leqslant k$ may depend on the independent variables $x^{i}$, the dependent variables $u^{i}$, and derivatives up to order $k$, while $\Delta_{i}^{\alpha}=0$ for $|\alpha|>k$.
For simplicity, we assume that the de Rham cohomology groups $H^{i}_{dR}(\mathcal{E})$ are trivial for $i>0$.
@@
The differential form $\omega|_{\mathcal{E}}$ represents a variational $1$-form of $\mathcal{E}$. This map from cosymmetries to variational $1$-forms is well-defined and surjective. The correspondence is an isomorphism if $\mathcal{E}$ is $\ell$-normal [19]. In this case, $\ker d_{1}^{\hskip 0.45206pt0,\hskip 0.60275ptn-1}=0$ and each conservation law $\xi\in E^{\hskip 0.3014pt0,\hskip 0.60275ptn-1}_{1}(\mathcal{E})$ is associated with a unique cosymmetry—namely, the one corresponding to $d_{1}\xi$. If $F=0$ is an Euler–Lagrange equation, then $l_{F}=l_{F}^{\,*}$ and cosymmetries of $\mathcal{E}$ can be regarded as its symmetries.
@@ 3 Cotangent coverings of $\ell$-normal systems
Remark 1.  The cotangent system $\mathcal{E}^{*}$ gives rise to the variational Schouten bracket and doesn’t depend on an $\ell$-normal embedding $\mathcal{E}\subset J^{\infty}(\pi)$ [3].
@@ 4 Hamiltonian operators
gives a characteristic of a conservation law of $\mathcal{E}$ for any $\psi_{1},\psi_{2}\in\widehat{P}(\pi)$. Then $l_{\mathcal{E}}^{\,*}\circ B|_{\mathcal{E}}=0$. Since $\mathcal{E}$ is $\ell$-normal, this condition implies $B|_{\mathcal{E}}=0$.
@@ 5 Invariant reduction mechanism
The invariant reduction of the $X$-invariant element represented by $\omega$ is the element of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq-1}(\mathcal{E}_{X})$ represented by $\vartheta|_{\mathcal{E}_{X}}$. We denote by $\mathcal{R}_{X}^{\hskip 0.3014ptp,\hskip 0.60275ptq}$ the homomorphism that evaluates the invariant reduction of $X$-invariant elements of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq}(\mathcal{E})$, provided it is well-defined. This is the case when $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq-1}(\mathcal{E})|_{\mathcal{E}_{X}}=~\!0$. If $E_{1}^{\hskip 0.3014pt0,\hskip 0.60275pt0}(\mathcal{E})|_{\mathcal{E}_{X}}=\mathbb{R}$, we also say that the invariant reduction of $X$-invariant elements of $E_{1}^{\hskip 0.3014pt0,\hskip 0.60275pt1}(\mathcal{E})$ is well-defined. The graded-commutative counterpart arises from the replacement of $(\mathcal{E},X)$ by $(\mathcal{E}^{*},\mathcal{X})$, where $\mathcal{X}$ is the lift of $X$. In this case, we additionally assume, without loss of generality, that $\omega$ and $\vartheta$ have the same internal degree.
@@
Theorem  1. Let $\mathcal{E}$ be an infinitely prolonged system of differential equations, and let $X$ be its evolutionary symmetry. Suppose that the invariant reduction is well-defined for $X$-invariant elements of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq}(\mathcal{E})$ and $E_{1}^{\hskip 0.3014ptp+1,\hskip 0.60275ptq}(\mathcal{E})$. Then on the space of $X$-invariant elements of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq}(\mathcal{E})$,
@@
Theorem  2. Suppose that $X$ and $X_{1}$ are commuting evolutionary symmetries of an infinitely prolonged system $\mathcal{E}$. If the invariant reduction is well-defined for $X$-invariant elements of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq}(\mathcal{E})$ and $E_{1}^{\hskip 0.3014ptp-1,\hskip 0.60275ptq}(\mathcal{E})$, then on the space of $X$-invariant elements of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq}(\mathcal{E})$,
$\mathcal{R}_{X}^{\hskip 0.3014ptp-1,\hskip 0.60275ptq}\circ{X_{1}}\lrcorner=-X_{1}|_{\mathcal{E}_{X}}\lrcorner\,\circ\mathcal{R}_{X}^{\hskip 0.3014ptp,\hskip 0.60275ptq}$
@@ 6 Cotangent coverings of finite-dimensional systems
Let $\pi_{\mathcal{S}}\colon\mathcal{S}^{n+N}\to M^{n}$ be a smooth, surjective submersion between finite-dimensional connected smooth manifolds. We assume that it is equipped with a structure of a differential equation, i.e., a flat connection [^7] It is a differential equation in a sense that is more general than the one adopted in Section 2.2. The connection canonically defines an embedding of $\mathcal{S}$ into the manifold $J^{1}(\pi_{\mathcal{S}})$ of $1$-jets of local sections of the submersion $\pi_{\mathcal{S}}$.. Denote by $D^{v}(\mathcal{S})$ the module of $\pi_{\mathcal{S}}$-vertical derivations on $\mathcal{S}$. The decomposition
@@
7 On finite-dimensional reductions
@@
In other words, the bivector maps the differential of the constant of motion $\mathcal{R}_{{X}}({\xi})$ to the symmetry ${X}_{\nabla(d_{1}\xi)}|_{\mathcal{E}_{{X}}}$. Then the relation between conservation laws and symmetries determined by $\nabla$ is, up to sign, inherited by the bivector corresponding to $\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})$. The reduction of (4.3) shows that for $X$-invariant conservation laws $\xi_{1},\xi_{2}\in E^{\hskip 0.3014pt0,\hskip 0.60275ptn-1}_{1}(\mathcal{E})$, the Poisson bracket of their reductions defined by $\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})$ coincides with
$-\mathcal{R}_{X}(\{\xi_{1},\xi_{2}\}_{\nabla})\,.$
@@
7.1 On reductions of (1+1)-dimensional evolution systems
Remark 6 shows that the existence of the isomorphism between $\mathfrak{T}^{*}[1]\mathcal{E}_{X}$ and $\mathcal{E}^{*}_{\mathcal{X}}$ considered above depends only on some non-degeneracy [^9] More precisely, the following operator must be a (well-defined) isomorphism of the $\mathcal{F}(\mathcal{E}_{X})$-module $D^{v}(\mathcal{E}_{X})$ and the $\mathcal{F}(\mathcal{E}_{X})$-module of degree-1 functions on $\mathcal{E}^{*}_{\mathcal{X}}$. It takes $w\in D^{v}(\mathcal{E}_{X})$, lifts it to $\mathcal{E}^{*}_{\mathcal{X}}$ in such a way that the lift preserves $\mathcal{R}_{\mathcal{X}}(\rho)$, and then contracts this lift with $\mathcal{R}_{\mathcal{X}}(\rho)$. Then it can be identified with $\mathbb{I}\in\mathrm{End}(D^{v}(\mathcal{E}_{X}))$. of the differential form $\mathcal{R}_{\mathcal{X}}(\rho)$. In this section, we prove the existence of such isomorphisms for some reductions of $(1+1)$-dimensional evolution systems and derive an explicit formula for the inherited Poisson bivectors.
@@ 7.1 On reductions of (1+1)-dimensional evolution systems
$u^{1}_{t}=f^{1}\,,\qquad\ldots\qquad u^{m}_{t}=f^{m}\,.$
Here $F^{i}=u^{i}_{t}-f^{i}$, while $f^{i}$ are functions of the independent variables $t$ and $x$, the dependent variables $u^{i}$, and their $x$-derivatives $u^{i}_{kx}$ up to some finite order. For $(1+1)$-dimensional evolution systems we also use the notation $u^{i}_{k}=u^{i}_{kx}$. As intrinsic coordinates on $\mathcal{E}$, we take $t$, $x$, and all variables of the form $u^{i}_{k}$, including $u^{i}_{0}=u^{i}$. Similarly, as intrinsic coordinates on $\mathcal{E}^{*}$, we use $t$, $x$, and all variables of the form $u^{i}_{k}$, $p_{i\,k}=p_{i\,kx}$.
@@
Let $X=E_{\varphi}|_{\mathcal{E}}$ be a symmetry of $\mathcal{E}$. The subsystem $\mathcal{E}_{X}\subset\mathcal{E}$, defined by the condition $X=0$, is given by the relations $D_{x}^{\hskip 0.3014ptk}(\varphi^{i})|_{\mathcal{E}}=0$. Without loss of generality, we assume that the components $\varphi^{1},\ldots,\varphi^{m}$ don’t depend on variables of the form $D_{\alpha}(u^{i}_{t})$. Then $E_{\varphi}(F)=l_{\varphi}(F)$ and the lift $\mathcal{X}$ has the characteristic $(\varphi,-l_{\varphi}^{\,*}(p))$. Suppose that for some integers $k_{1}$, …, $k_{m}\geqslant-1$, the determinant
@@
respectively. Equating the restrictions of these coefficients to the variables $\varpi_{i\,k_{i}}$, $\varpi_{i\,k_{i}-1}$, $\varpi_{i\,k_{i}-2}$, …, $\varpi_{i\,0}$, respectively, we get the desired isomorphism of $\mathfrak{T}^{*}[1]\mathcal{E}_{X}$ and $\mathcal{E}^{*}_{\mathcal{X}}$. Indeed, the following relations can be solved for the variables $p_{j}$
@@
Here the restriction to $\mathcal{E}^{*}_{\mathcal{X}}$ can be replaced with the restriction to $\mathcal{E}_{X}$ or simply to $\varphi=0$. Then, at the next step, the variables $p_{j\,1}$ can be determined from the expressions for $\varpi_{i\,k_{i}-1}$, and so on. Therefore, if $\nabla\colon\widehat{P}(\mathcal{E})\to\varkappa(\mathcal{E})$ is an $X$-invariant Hamiltonian operator, then there is a Poisson bracket on $E^{\hskip 0.3014pt0,\hskip 0.60275pt0}_{1}(\mathcal{E}_{X})$, inherited from $\{\cdot,\cdot\}_{\nabla}$ up to sign. It is defined by $\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})$. We obtain
@@
$\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})\simeq\dfrac{1}{2}\Big(\hat{\nabla}^{ij\,k}w_{j\,k}\wedge\partial_{u^{i}_{0}}+\mathcal{L}_{\widetilde{D}_{x}}(\hat{\nabla}^{ij\,k}w_{j\,k})\wedge\partial_{u^{i}_{1}}+\ldots+\mathcal{L}_{\widetilde{D}_{x}}^{\,k_{i}}(\hat{\nabla}^{ij\,k}w_{j\,k})\wedge\partial_{u^{i}_{k_{i}}}\Big),$
@@
$-\dfrac{1}{2}\Big(\hat{\nabla}^{ij\,k}w_{j\,k}\wedge\partial_{u^{i}_{0}}+\mathcal{L}_{\widetilde{D}_{x}}(\hat{\nabla}^{ij\,k}w_{j\,k})\wedge\partial_{u^{i}_{1}}+\ldots+\mathcal{L}_{\widetilde{D}_{x}}^{\,k_{i}}(\hat{\nabla}^{ij\,k}w_{j\,k})\wedge\partial_{u^{i}_{k_{i}}}\Big),\qquad\hat{\nabla}^{ij\,k}=\nabla^{ij\,k}|_{\mathcal{E}_{X}},$
@@
Remark 15.  If $X$ is a Hamiltonian symmetry corresponding to $\xi\in E^{\hskip 0.3014pt0,\hskip 0.60275pt1}_{1}(\mathcal{E})$, i.e., $X=X_{\nabla(d_{1}\xi)}$, then $\mathcal{R}_{X}(\xi)$ is a Casimir function for the bivector corresponding to $\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})$. As with reductions of other conservation laws, it is nontrivial if and only if the cosymmetry of $\xi$ doesn’t vanish on $\mathcal{E}_{X}$ (see Remark 7 in [4]) provided, for example, that the assumptions of Theorem 7.1 are satisfied.

### v3

@@ Abstract
We show that, under suitable conditions, finite-dimensional systems describing invariant solutions of partial differential equations (PDEs) inherit local Hamiltonian operators through the mechanism of invariant reduction, which applies uniformly to point, contact, and higher symmetries. The inherited operators endow the reduced systems with Poisson bivectors that relate constants of motion to symmetries. Applying the same mechanism to invariant conservation laws, we further show that the induced Poisson brackets agree with those of the original systems, up to sign. The results are illustrated by two examples in which the inherited Poisson brackets and inherited constants of motion yield integrability of the reduced systems. Crucially, the construction is independent of the choice of an $\ell$-normal inclusion of a PDE system into jet spaces.
@@ 1 Introduction
Our mechanism of invariant reduction is related to methods introduced in [5] and [6]. As formulated in those works, these methods are not directly applicable to higher symmetries, but they do allow one to consider multi-reduction in a more general setting. Moreover, following the ideas of this paper, a slight extension of the method in [5] to the reduction of the invariant part of the zero page of Vinogradov’s $\mathcal{C}$-spectral sequence for a given equation may allow Poisson bracket reductions under suitable, in some cases non-solvable, multi-dimensional Lie algebras of point or contact symmetries.
@@
Aside from this, there is a substantial literature on related questions concerning finite-dimensional symmetry reductions and the inheritance of Poisson brackets for evolution systems (see, e.g., [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]). Our approach is distinguished by its generality and systematic nature: assuming suitable regularity conditions, it applies to $\ell$-normal systems of partial differential equations, with $(1+1)$-dimensional evolution systems included as a particular case. We also emphasize that our reduction mechanism is global and geometric. In particular, it does not rely on local coordinates, auxiliary variables for reduced systems, or the existence of Lax pairs. Furthermore, it is canonical in the sense that it is independent of the choice of an $\ell$-normal inclusion of a PDE system into jet spaces. While some aspects of symmetry reductions and Poisson bracket inheritance have been studied previously, the generality and formulation of our approach – including Theorem 7.1 – appear to be novel.
@@ 2.1 Jets
Here $\langle\cdot,\cdot\rangle$ denotes the natural pairing between a module and its adjoint. Note that $\mathrm{E}(L)$ can be identified with the differential form that takes the following form in adapted coordinates:
@@ 2.2 Differential equations
Regularity assumptions. We say that $\mathcal{E}\subset J^{\infty}(\pi)$ is regular if the following condition holds. A function $f\in\mathcal{F}(\pi)$ vanishes on $\mathcal{E}$ if and only if there is a differential operator $\Delta\colon P(\pi)\to\mathcal{F}(\pi)$ of the form $\Delta_{i}^{\alpha}D_{\alpha}$ ($\mathcal{C}$-differential operator or total differential operator) such that $f=\Delta(F)$. Here, for some integer $k$, the components $\Delta_{i}^{\alpha}$ with $|\alpha|\leqslant k$ may depend on the independent variables $x^{i}$, the dependent variables $u^{i}$, and derivatives up to order $k$, while $\Delta_{i}^{\alpha}=0$ for $|\alpha|>k$. We consider only regular systems. For simplicity, we assume that the de Rham cohomology groups $H^{i}_{dR}$ of all systems under consideration are trivial for $i>0$.
@@
The differential form $\omega|_{\mathcal{E}}$ represents a variational $1$-form of $\mathcal{E}$. This map from cosymmetries to variational $1$-forms is well-defined and surjective. The correspondence is an isomorphism if $\mathcal{E}$ is $\ell$-normal [20]. In this case, $\ker d_{1}^{\hskip 0.45206pt0,\hskip 0.60275ptn-1}=0$ and each conservation law $\xi\in E^{\hskip 0.3014pt0,\hskip 0.60275ptn-1}_{1}(\mathcal{E})$ is associated with a unique cosymmetry – namely, the one corresponding to $d_{1}\xi$. If $F=0$ is an Euler–Lagrange equation, then $l_{F}=l_{F}^{\,*}$ and cosymmetries of $\mathcal{E}$ can be regarded as its symmetries.
@@ 3 Cotangent coverings of $\ell$-normal systems
Remark 1.  The cotangent system $\mathcal{E}^{*}$ gives rise to the variational Schouten bracket and, along with its canonical variational $1$-form, is independent of the choice of $\ell$-normal inclusion $\mathcal{E}\subset J^{\infty}(\pi)$ [3].
@@ 4 Hamiltonian operators
yields a characteristic of a conservation law of $\mathcal{E}$ for any $\psi_{1},\psi_{2}\in\widehat{P}(\pi)$. Then $l_{\mathcal{E}}^{\,*}\circ B|_{\mathcal{E}}=0$. Since $\mathcal{E}$ is $\ell$-normal, this condition implies $B|_{\mathcal{E}}=0$.
@@ 4.2 Variational bivectors as odd symmetries
Remark 3.  The identification of the mutually corresponding variational bivectors arising from different $\ell$-normal inclusions of $\mathcal{E}$ into jet spaces (assuming that $\mathcal{E}^{*}$ is also $\ell$-normal) allows us to conclude that the symmetry $s_{\nabla}$ and the subsequent reduction mechanism are independent of the choice of such an inclusion, thereby ensuring the canonical nature of our approach.
@@ 5 Invariant reduction mechanism
The invariant reduction of the $X$-invariant element represented by $\omega$ is the element of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq-1}(\mathcal{E}_{X})$ represented by $\vartheta|_{\mathcal{E}_{X}}$. We denote by $\mathcal{R}_{X}^{\hskip 0.3014ptp,\hskip 0.60275ptq}$ the homomorphism that evaluates the invariant reduction of $X$-invariant elements of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq}(\mathcal{E})$, provided it is well-defined. This is the case when $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq-1}(\mathcal{E})|_{\mathcal{E}_{X}}=~\!0$. If $E_{1}^{\hskip 0.3014pt0,\hskip 0.60275pt0}(\mathcal{E})|_{\mathcal{E}_{X}}\subset H^{0}_{dR}(\mathcal{E}_{X})$, we also say that the invariant reduction of $X$-invariant elements of $E_{1}^{\hskip 0.3014pt0,\hskip 0.60275pt1}(\mathcal{E})$ is well-defined, as we consider their reductions up to additive locally constant functions on $\mathcal{E}_{X}$. The graded-commutative counterpart arises from the replacement of $(\mathcal{E},X)$ by $(\mathcal{E}^{*},\mathcal{X})$, where $\mathcal{X}$ is the lift of $X$. In this case, we additionally assume, without loss of generality, that $\omega$ and $\vartheta$ have the same internal degree.
@@
Theorem  1. Let $\mathcal{E}$ be an infinitely prolonged system of differential equations, and let $X=E_{\varphi}|_{\mathcal{E}}$ be its symmetry. Suppose that the invariant reduction is well-defined for $X$-invariant elements of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq}(\mathcal{E})$ and $E_{1}^{\hskip 0.3014ptp+1,\hskip 0.60275ptq}(\mathcal{E})$. Then on the $X$-invariant subspace of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq}(\mathcal{E})$,
@@
Theorem  2. Suppose that $X=E_{\varphi}|_{\mathcal{E}}$, $X_{1}=E_{\varphi_{1}}|_{\mathcal{E}}$ are commuting symmetries of an infinitely prolonged system $\mathcal{E}$. If the invariant reduction is well-defined for $X$-invariant elements of $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq}(\mathcal{E})$ and $E_{1}^{\hskip 0.3014ptp-1,\hskip 0.60275ptq}(\mathcal{E})$, then on the $X$-invariant subspace of  [^7] If $(p;q)=(1;1)$, the relation (5.1) is understood up to an additive locally constant function on $\mathcal{E}_{X}$. $E_{1}^{\hskip 0.3014ptp,\hskip 0.60275ptq}(\mathcal{E})$,
$\mathcal{R}_{X}^{\hskip 0.3014ptp-1,\hskip 0.60275ptq}\circ{X_{1}}\lrcorner=-X_{1}|_{\mathcal{E}_{X}}\lrcorner\,\circ\mathcal{R}_{X}^{\hskip 0.3014ptp,\hskip 0.60275ptq}$ (5.1)
@@ 6 Cotangent coverings of finite-dimensional systems
Let $\pi_{\mathcal{S}}\colon\mathcal{S}^{n+N}\to M^{n}$ be a smooth, surjective submersion between finite-dimensional connected smooth manifolds. We assume that it is equipped with a structure of a differential equation, i.e., a flat connection [^8] It is a differential equation in a sense that is more general than the one adopted in Section 2.2. The connection canonically defines an embedding of $\mathcal{S}$ into the manifold $J^{1}(\pi_{\mathcal{S}})$ of $1$-jets of local sections of the submersion $\pi_{\mathcal{S}}$. We do not require the connection to have global path lifting.. Denote by $D^{v}(\mathcal{S})$ the module of $\pi_{\mathcal{S}}$-vertical derivations on $\mathcal{S}$. The decomposition
@@
7 Finite-dimensional reductions
@@
In other words, the bivector maps the differential of $\mathcal{R}_{{X}}({\xi})$ to the symmetry ${X}_{\nabla(d_{1}\xi)}|_{\mathcal{E}_{{X}}}$. Then the relation between conservation laws and symmetries determined by $\nabla$ is, up to sign, inherited by the bivector corresponding to $\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})$. The reduction of (4.3) shows that for $X$-invariant conservation laws $\xi_{1},\xi_{2}\in E^{\hskip 0.3014pt0,\hskip 0.60275ptn-1}_{1}(\mathcal{E})$, the Poisson bracket of their reductions defined by $\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})$ represents [^10] Since we consider the case $n=2$ (and $H^{0}_{dR}(\mathcal{E}_{X})=\mathbb{R}$), the reduction (7.1) is defined up to an additive real number. In particular, if two $X$-invariant conservation laws of $\mathcal{E}$ commute with respect to $\{\cdot,\cdot\}_{\nabla}$, the inherited Poisson bracket of their reductions is a real number.
$-\mathcal{R}_{X}(\{\xi_{1},\xi_{2}\}_{\nabla})\,.$ (7.1)
@@
7.1 Reductions of (1+1)-dimensional evolution systems
Remark 6 shows that the existence of the isomorphism between $\mathfrak{T}^{*}[1]\mathcal{E}_{X}$ and $\mathcal{E}^{*}_{\mathcal{X}}$ considered above depends only on some non-degeneracy of the differential form $\mathcal{R}_{\mathcal{X}}(\rho)$. In this section, we prove the existence of such isomorphisms for some reductions of $(1+1)$-dimensional evolution systems and derive an explicit formula for the inherited Poisson bivectors.
@@ 7.1 Reductions of (1+1)-dimensional evolution systems
$u^{1}_{t}=f^{1}\,,\qquad\ldots\,,\qquad u^{m}_{t}=f^{m}\,.$
Here $\zeta=\pi$. In adapted local coordinates, $F^{i}=u^{i}_{t}-f^{i}$, while $f^{i}$ are functions of the independent variables $t$ and $x$, the dependent variables $u^{i}$, and their $x$-derivatives $u^{i}_{kx}$ up to some finite order. For $(1+1)$-dimensional evolution systems we also use the notation $u^{i}_{k}=u^{i}_{kx}$. As intrinsic coordinates on $\mathcal{E}$, we take $t$, $x$, and all variables of the form $u^{i}_{k}$, including $u^{i}_{0}=u^{i}$. Similarly, as intrinsic coordinates on $\mathcal{E}^{*}$, we use $t$, $x$, and all variables of the form $u^{i}_{k}$, $p_{i\,k}=p_{i\,kx}$.
@@
Let $X=E_{\varphi}|_{\mathcal{E}}$ be a symmetry of $\mathcal{E}$. The subsystem $\mathcal{E}_{X}\subset\mathcal{E}$, defined by the condition $X=0$, is given by the relations $D_{x}^{\hskip 0.3014ptk}(\varphi^{i})|_{\mathcal{E}}=0$. Without loss of generality, we assume that the components $\varphi^{1},\ldots,\varphi^{m}$ do not depend on variables of the form $D_{\alpha}(u^{i}_{t})$. Then $E_{\varphi}(F)=l_{\varphi}(F)$ and the lift $\mathcal{X}$ has the characteristic $(\varphi,-l_{\varphi}^{\,*}(p))$. Suppose that for some integers $k_{1}$, …, $k_{m}\geqslant-1$, the determinant
@@
respectively. Equating the restrictions of these coefficients to the variables $\varpi_{i\,k_{i}}$, $\varpi_{i\,k_{i}-1}$, $\varpi_{i\,k_{i}-2}$, …, $\varpi_{i\,0}$, respectively, we obtain the desired isomorphism of $\mathfrak{T}^{*}[1]\mathcal{E}_{X}$ and $\mathcal{E}^{*}_{\mathcal{X}}$. Indeed, the following relations can be solved for the variables $p_{j}$
@@
Here the restriction to $\mathcal{E}^{*}_{\mathcal{X}}$ can be replaced with the restriction to $\mathcal{E}_{X}$ or simply to $\varphi=0$. Then, at the next step, the variables $p_{j\,1}$ can be determined from the expressions for $\varpi_{i\,k_{i}-1}$, and so on. One can see that, in local coordinates under consideration, there exists a unique lift of a vertical derivation from $\mathcal{E}_{X}$ to $\mathcal{E}_{\mathcal{X}}^{*}$ preserving $\mathcal{R}_{\mathcal{X}}(\rho)$. Its uniqueness ensures that the lift is well-defined globally. As it follows from the coordinate expression for $\mathcal{R}_{\mathcal{X}}(\rho)$, the operator that contracts such lifts with $\mathcal{R}_{\mathcal{X}}(\rho)$ establishes an isomorphism between the $\mathcal{F}(\mathcal{E}_{X})$-modules of vertical derivations on $\mathcal{E}_{X}$ and degree-$1$ functions on $\mathcal{E}^{*}_{\mathcal{X}}$. Therefore, it yields the identification of $\mathfrak{T}^{*}[1]\mathcal{E}_{X}$ and $\mathcal{E}^{*}_{\mathcal{X}}$ relating the corresponding $1$-form $\rho_{c}$ to $\mathcal{R}_{\mathcal{X}}(\rho)$. Its coordinate form is given by the expressions for $\varpi_{i\,k_{i}}$, $\varpi_{i\,k_{i}-1}$, $\ldots,$ $\varpi_{i\,0}$.
Thus, if $\nabla\colon\widehat{P}(\mathcal{E})\to\varkappa(\mathcal{E})$ is an $X$-invariant Hamiltonian operator, then there is a Poisson bracket on $E^{\hskip 0.3014pt0,\hskip 0.60275pt0}_{1}(\mathcal{E}_{X})$, inherited from $\{\cdot,\cdot\}_{\nabla}$ up to sign. It is defined by $\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})$. We obtain
@@
$\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})\simeq\dfrac{1}{2}\Big(\hat{\nabla}^{ij\,k}w_{j\,k}\wedge\partial_{u^{i}_{0}}+\mathcal{L}_{\widetilde{D}_{x}}(\hat{\nabla}^{ij\,k}w_{j\,k})\wedge\partial_{u^{i}_{1}}+\mathcal{L}_{\widetilde{D}_{x}}^{\,2}(\hat{\nabla}^{ij\,k}w_{j\,k})\wedge\partial_{u^{i}_{2}}+\ldots\Big),$
@@
$-\dfrac{1}{2}\sum_{r}\mathcal{L}_{\widetilde{D}_{x}}^{\,r}(\hat{\nabla}^{ij\,k}w_{j\,k})\wedge\partial_{u^{i}_{r}}\,,\qquad\hat{\nabla}^{ij\,k}=\nabla^{ij\,k}|_{\mathcal{E}_{X}},$
@@
Remark 16.  If $X$ is a Hamiltonian symmetry corresponding to $\xi\in E^{\hskip 0.3014pt0,\hskip 0.60275pt1}_{1}(\mathcal{E})$, i.e., $X=X_{\nabla(d_{1}\xi)}$, then $\mathcal{R}_{X}(\xi)$ is a Casimir function [^11] The bivector corresponding to $\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})$ relates $\mathcal{R}_{X}(\xi)$ to the symmetry $X|_{\mathcal{E}_{X}}=0$. for the bivector corresponding to $\mathcal{R}_{\mathcal{X}}(\mathcal{H}_{\nabla})$. As with reductions of other conservation laws, it is nontrivial if and only if the cosymmetry of $\xi$ does not vanish on $\mathcal{E}_{X}$ (see Remark 7 in [4]) provided, for example, that the assumptions of Theorem 7.1 are satisfied.